# TURA: Tool-Augmented Unified Retrieval Agent for AI Search

### v1

@@ Abstract.
The advent of Large Language Models (LLMs) is transforming search engines into conversational AI search products, primarily using Retrieval-Augmented Generation (RAG) on web corpora. However, this paradigm has significant industrial limitations. Traditional RAG approaches struggle with real-time needs and structured queries that require accessing dynamically generated content like ticket availability or inventory. Limited to indexing static pages, search engines cannot perform the interactive queries needed for such time-sensitive data. Academic research has focused on optimizing RAG for static content, overlooking complex intents and the need for dynamic sources like databases and real-time APIs. To bridge this gap, we introduce TURA (Tool-Augmented Unified Retrieval Agent for AI Search), a novel three-stage framework that combines RAG with agentic tool-use to access both static content and dynamic, real-time information. TURA has three key components: an Intent-Aware Retrieval module to decompose queries and retrieve information sources encapsulated as Model Context Protocol (MCP) Servers, a DAG-based Task Planner that models task dependencies as a Directed Acyclic Graph (DAG) for optimal parallel execution, and a lightweight Distilled Agent Executor for efficient tool calling. TURA is the first architecture to systematically bridge the gap between static RAG and dynamic information sources for a world-class AI search product. Serving tens of millions of users, it leverages an agentic framework to deliver robust, real-time answers while meeting the low-latency demands of a large-scale industrial system.
Conversational Search; Large Language Models; Information Retrieval
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Question answering
@@ 1. Introduction
Figure 1. Demonstration of TURA’s agentic capabilities. Given a query on July 31, 2025: (a) TURA autonomously utilizes a tool by calling Ctrip’s API to successfully look up ticket information. (b) This contrasts with a traditional RAG-based AI search, which can only retrieve information from static webpages and is incapable of performing the required action.
[…]
@@
Traditional web search engines, built on foundational algorithms like PageRank (Page et al., 1999) and the early Google architecture (Brin and Page, 1998), have long relied on the “ten blue links” paradigm. While these systems excel at retrieving and ranking information from vast corpora of unstructured web pages, they have historically struggled with queries demanding structured, real-time, or transactional information, such as flight availability or weather forecasts. To address these limitations, search engines introduced manually curated components like Google’s OneBox (Elia, 2017) or Baidu’s Aladdin platform (Zhang and Duan, 2012), which present information from specific, trusted data sources in dedicated formats. However, this approach proves fundamentally limited by its reliance on brittle, hand-crafted integrations that are difficult to scale.
The advent of Large Language Models (LLMs) has catalyzed a fundamental paradigm shift in information access (Chowdhury and Chowdhury, 2024), transforming search from keyword-based retrieval to conversational, answer-centric systems. The dominant architecture in this new era is Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Guu et al., 2020), which grounds LLM responses in factual knowledge retrieved from web-scale corpora. Commercial deployments like Perplexity AI, ChatGPT’s web search integration, and Google’s AI Overviews demonstrate the practical viability of this approach (Xiong et al., 2024).
@@
To bridge this critical gap between retrieving from static corpora and interacting with dynamic data sources, we propose TURA (Tool-Augmented Unified Retrieval Agent for AI Search). TURA is a novel three-stage agentic framework that enhances LLMs with the ability to use external tools, moving beyond passive document retrieval to active, real-time data acquisition. The system’s design follows the ReAct framework (Yao et al., 2023) of interleaving reasoning and acting, and incorporates advanced planning capabilities inspired by adaptive decomposition methods (Prasad et al., 2023). For efficient execution, TURA implements DAG-based task decomposition for parallel tool calls (Kim et al., 2024), addressing latency challenges while maintaining the complex reasoning capabilities demonstrated in large-scale API integration frameworks (Qin et al., 2023). TURA’s architecture leverages standardized tool interfaces following the Model Context Protocol (MCP) specification (Anthropic, 2024), enabling seamless integration with diverse information sources through a unified framework.
@@
TURA’s three-stage architecture comprises the following components: (1) Intent-Aware Tool Retrieval: This module decomposes complex user queries into atomic sub-intents and performs dense retrieval over a semantically-indexed catalogue of available tools, encapsulating both static document collections and dynamic APIs, to identify relevant candidates for each sub-intent. (2) DAG-based Task Planning: This module constructs optimal and parallelizable execution plans by modeling sub-tasks and their data dependencies as a Directed Acyclic Graph (DAG), enabling the orchestration of complex, multi-hop reasoning chains across multiple tool calls. (3) Distilled Agent Executor: To address critical inference latency barriers in production settings, we developed a lightweight yet highly capable agent fine-tuned on curated expert trajectories using a novel mixed-rationale distillation technique, achieving comparable fidelity to proprietary LLMs at a fraction of the computational cost and latency.
Since May 2025, TURA has been fully deployed, successfully serving tens of millions of users. TURA demonstrably expands the capabilities of AI search, providing accurate, real-time answers for a broad spectrum of queries, particularly those involving dynamic, transactional, or non-web data, that were previously intractable for conventional RAG-based systems.
Our contributions are threefold:
@@
  • We propose TURA, a novel and systematic agentic architecture that effectively integrates diverse, dynamic and non-web information sources into AI search via tool calling, directly addressing the static-world limitations of traditional RAG systems.
@@
  • We introduce a cohesive framework of synergistic techniques including intent-aware tool retrieval, DAG-based planning, and a latency-optimized distilled agent that collectively solve key industrial challenges of tool selection, planning, and efficient execution.
  • We present the first large-scale industrial validation of a tool-augmented agentic search system, demonstrating its viability and effectiveness in a world-class AI search product. This provides a production-proven blueprint for the next generation of AI search.
@@ 2.1. Retrieval-Augmented Generation
To mitigate LLM hallucinations and improve factual accuracy (Guu et al., 2020; Fan et al., 2024), RAG systems ground the generation process in external knowledge. These systems typically employ a ”retrieve-then-generate” paradigm, with a significant body of research focused on optimizing both retrieval (Ma et al., 2023; Tao et al., 2024; Lee et al., 2025) and generation (Liu et al., 2023; Wang et al., 2024a). Recent advances in RAG have progressed from static retrieve-then-generate pipelines to more dynamic and adaptive frameworks (Wang et al., 2024b; Fan et al., 2024). For example, Self-RAG introduces ”reflection tokens” that enable the model to decide on-demand whether retrieval is necessary, avoiding the inefficiency of indiscriminate retrieval for every query (Asai et al., 2024). In a similar vein, Active Retrieval Augmented Generation proposes an iterative retrieval process that occurs throughout generation, allowing the model to gather information as needed. Meanwhile, R²AG focuses on bridging the semantic gap between the retriever and the generator by using more nuanced retrieval features (Ye et al., 2024). Despite these innovations, the practical application of RAG in commercial systems like Perplexity AI and Google AI Overviews has highlighted ongoing challenges in maintaining content quality and robust contextual understanding (Williams, 2024). A fundamental and pervasive limitation of existing systems is their foundational reliance on rigid, predefined workflows. This inherent inflexibility means they are ill-equipped to dynamically adapt to the demands of complex, multi-faceted queries, which by their very nature necessitate a seamless and intelligent synthesis of information drawn from a wide array of diverse sources and the orchestrated application of various computational tools.
@@ 2.2. Tool-Augmented Agents
The foundational ReAct framework established a paradigm of interleaving reasoning and acting, where the model generates both thought processes and subsequent actions to interact with its environment. This synergy was shown to significantly improve task performance and interpretability. Building on this, Toolformer demonstrated how models could teach themselves to use tools in a self-supervised manner (Schick et al., 2023). The scale of tool integration was dramatically expanded by ToolLLM, which enabled models to leverage thousands of real-world APIs through the use of depth-first search-based decision trees (Qin et al., 2023). Recent work has also focused on optimizing the efficiency of tool use. LLMCompiler, for instance, introduced a Directed Acyclic Graph (DAG)-based approach for parallel tool calling, achieving a 3.6x speedup (Kim et al., 2024). Agent Q further improved success rates by combining Monte Carlo Tree Search with a self-critique mechanism (Putta et al., 2024).
Despite these advances, current systems face critical limitations: (1) their workflows are often static and cannot adapt to the complexity of a given query; (2) they struggle with the semantic integration of heterogeneous tools and information sources; and (3) they suffer from coordination inefficiencies, as RAG and tool-augmented systems typically operate in isolation. TURA is designed to address these issues by proposing a unified architecture that dynamically coordinates retrieval decisions, tool selection, and response generation based on the specific characteristics and intent of the user’s query.
@@ 3. Problem Definition
We are given a user’s natural language query, $q$, and a collection of $N$ heterogeneous MCP Servers, $M=\{M_{1},M_{2},\dots,M_{N}\}$. These servers function as external, specialized tools that provide access to information and capabilities not inherently present in a language model.
The core problem is to answer the query $q$ by dynamically composing the functionalities of these servers. This involves selecting the right servers and coordinating their execution to produce a final, synthesized answer, $A$. We refer to any such specific composition of server calls as an execution strategy, denoted by $\pi$.
The effectiveness of a strategy is determined by a fundamental trade-off. We aim to maximize the quality of the final answer, $Q(A)$, which depends on the chosen strategy, while simultaneously adhering to a strict latency constraint, $L(\pi)$. Our objective is to find the optimal strategy, $\pi^{*}$, that yields the best possible answer within a given time budget. This is formalized as the following constrained optimization problem:
(1) $\pi^{*}=\operatornamewithlimits{arg\,max}_{\pi}Q(A(\pi))\quad\text{subject to}\quad L(\pi)\leq\tau_{\text{max}}$
where $\tau_{\text{max}}$ represents the maximum permissible latency. This formulation encapsulates the challenge of leveraging external tools effectively under real-world performance requirements.
Figure 2. TURA Framework Overview. The framework consists of three stages: Intent-Aware MCP Server Retrieval, DAG-based Task Planner, and Distilled Agent Executor. Example shows processing a Beijing travel query.
[…]
@@
4. Method
Fig. 2 illustrates the architecture of the TURA, which consists of three key modules: an Intent-Aware MCP Server Retrieval Module, a DAG-based Task Planner Module, and a latency-optimized Distilled Agent Executor Module.
@@ 4.1.1. LLM-based Multi-Intent Query Decomposition
User queries are often underspecified and multi-faceted. A monolithic query may contain multiple, logically distinct information needs. To handle this, we employ a powerful LLM, $f_{\text{LLM-de}}$, configured with a specific prompt (see Appendix A.2) to function as a query decomposer. The LLM is instructed to parse the raw query $q$ and transform it into a structured set of atomic sub-queries, $\mathcal{SQ}=\{sq_{1},sq_{2},\dots,sq_{k}\}$.
(2) $\mathcal{SQ}=f_{\text{LLM-de}}(q)$
Each sub-query $sq_{j}\in\mathcal{SQ}$ is designed to be unambiguous and correspond to a single semantic intent (e.g., ”find the city where the Forbidden City is located.”, ”get weather forecast for a given city”). This decomposition transforms an ambiguous problem into a set of well-defined, tractable sub-problems. This strategy is inspired by recent advances in task decomposition for complex reasoning (Zhou et al., 2022).

### v2

@@ Abstract.
The advent of Large Language Models (LLMs) is transforming search engines into conversational AI search products, primarily using Retrieval-Augmented Generation (RAG) on web corpora. However, this paradigm has significant industrial limitations. Traditional RAG approaches struggle with real-time needs and structured queries that require accessing dynamically generated content like ticket availability or inventory. Limited to indexing static pages, search engines cannot perform the interactive queries needed for such time-sensitive data. Academic research has focused on optimizing RAG for static content, overlooking complex intents and the need for dynamic sources like databases and real-time APIs. To bridge this gap, we introduce TURA (Tool-Augmented Unified Retrieval Agent for AI Search), a novel three-stage framework that combines RAG with agentic tool-use to access both static content and dynamic, real-time information. TURA has three key components: an Intent-Aware Retrieval module to decompose queries and retrieve information sources encapsulated as Model Context Protocol (MCP) Servers, a DAG-based Task Planner that models task dependencies as a Directed Acyclic Graph (DAG) for optimal parallel execution, and a lightweight Distilled Agent Executor for efficient tool calling. Moreover, we release a comprehensive benchmark derived from anonymized Baidu Search production logs with expert annotations, covering diverse static–dynamic query scenarios and standardized evaluation protocols. TURA is the first architecture to systematically bridge the gap between static RAG and dynamic information sources for a world-class AI search product. Serving tens of millions of users, it leverages an agentic framework to deliver robust, real-time answers while meeting the low-latency demands of a large-scale industrial system.
Large Language Models, Retrieval-Augmented Generation, Multi-Agent System, Question Answering, AI Search
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXsubmissionid: 123-A56-BU3copyright: nonefootnotetext: Equal contribution.
@@ 1. Introduction
Figure 1. Demonstration of TURA’s agentic capabilities. Given a query on July 31, 2025: (a) TURA autonomously utilizes a tool by calling Ctrip’s API to successfully look up ticket information and complete the required action. (b) This contrasts with a traditional RAG-based AI search, which can only retrieve information from static webpages and is unable to perform the required action.
@@
Traditional web search engines, built on foundational algorithms like PageRank (Page et al., 1999) and the early Google architecture (Brin and Page, 1998), have long relied on the “ten blue links” paradigm (Li et al., 2025b). While these systems excel at retrieving and ranking information from vast corpora of unstructured web pages, they have historically struggled with queries demanding structured, real-time, or transactional information, such as flight availability or weather forecasts. To address these limitations, search engines introduced manually curated components like Google’s OneBox (Elia, 2017) or Baidu’s Aladdin platform (Zhang and Duan, 2012; Li et al., 2023), which present information from specific, trusted data sources in dedicated formats. However, these approaches prove fundamentally limited by their reliance on brittle, hand-crafted integrations that are difficult to scale.
The advent of Large Language Models (LLMs) has catalyzed a fundamental paradigm shift in information access (Chowdhury and Chowdhury, 2024), transforming search from keyword-based retrieval to conversational, answer-centric systems. The dominant architecture in this new era is Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Guu et al., 2020), which grounds LLM responses in factual knowledge retrieved from web-scale corpora. Commercial deployments like Perplexity AI, ChatGPT’s web search integration, and Google’s AI Overviews demonstrate the practical viability of this approach (Xiong et al., 2024; Li et al., 2025a).
@@
To bridge this critical gap between retrieving from static corpora and interacting with dynamic data sources, we propose TURA (Tool-Augmented Unified Retrieval Agent for AI Search), which is a novel three-stage agentic framework that enhances LLMs with the ability to use external tools, moving beyond passive document retrieval to active, real-time data acquisition. The system’s design follows the ReAct framework (Yao et al., 2023), which interleaves reasoning and acting, and incorporates advanced planning capabilities inspired by adaptive decomposition methods (Prasad et al., 2023). For efficient execution, TURA implements DAG-based task decomposition for parallel tool calls (Kim et al., 2024), addressing latency challenges while maintaining the complex reasoning capabilities demonstrated in large-scale API integration frameworks (Qin et al., 2023). TURA’s architecture leverages standardized tool interfaces following the Model Context Protocol (MCP) specification (Anthropic, 2024), enabling seamless integration with diverse information sources through a unified framework. Specifically, TURA’s three-stage architecture comprises the following components: (1) Intent-Aware Tool Retrieval: This module decomposes complex user queries into atomic sub-intents and performs dense retrieval over a semantically-indexed catalogue of available tools, encapsulating both static document collections and dynamic APIs, to identify relevant candidates for each sub-intent. (2) DAG-based Task Planning: This module constructs optimal and parallelizable execution plans by modeling sub-tasks and their data dependencies as a Directed Acyclic Graph (DAG), enabling the orchestration of complex, multi-hop reasoning chains across multiple tool calls. (3) Distilled Agent Executor: To address critical inference latency barriers in production settings, TURA leverages a lightweight yet highly capable agent fine-tuned on curated expert trajectories using a novel mixed-rationale distillation technique, achieving comparable fidelity to proprietary LLMs at a fraction of the computational cost and latency. We conduct comprehensive offline experiments to rigorously evaluate the performance and efficiency of the proposed TURA framework. Offline experimental results reveal that TURA substantially improves answer accuracy and faithfulness compared with strong baselines, while comprehensive ablation studies confirm the indispensable role of each module. Moreover, since May 2025, TURA has been fully deployed, successfully serving tens of millions of users. TURA demonstrably expands the capabilities of AI search, providing accurate, real-time answers for a broad spectrum of queries, particularly those involving dynamic, transactional, or non-web data, that were previously intractable for conventional RAG-based systems. In summary, the contributions of this work could be categorized as follows:
  • We formulate AI search with heterogeneous information sources as a unified retrieval-and-execution problem over MCP servers, covering both static indexed corpora and dynamic transactional APIs, and optimize answer quality under strict latency constraints.
@@
  • We propose TURA, a novel and systematic agentic architecture that effectively integrates diverse, dynamic, and non-web information sources into AI search via tool calling, directly addressing the static-world limitations of traditional RAG systems.
@@
  • We introduce a cohesive framework of synergistic techniques, including intent-aware tool retrieval, DAG-based task planning, and a latency-optimized distilled agent that collectively solve key industrial challenges of tool selection, task planning, and efficient execution.
  • We present the first large-scale industrial validation of a tool-augmented agentic search system, demonstrating its viability and effectiveness in a world-class AI search product. In addition, we release MCP-Bench, a comprehensive benchmark constructed from anonymized production logs with expert annotations, providing standardized evaluation protocols and diverse static–dynamic query scenarios to facilitate reproducible research on tool-augmented AI search.
@@ 2.1. Retrieval-Augmented Generation
To mitigate LLM hallucinations and improve factual accuracy (Guu et al., 2020; Fan et al., 2024), RAG systems ground the generation process in external knowledge. These systems typically employ a ”retrieve-then-generate” paradigm, with a significant body of research focused on optimizing both retrieval (Ma et al., 2023; Tao et al., 2024; Lee et al., 2025) and generation (Liu et al., 2023; Wang et al., 2024a). Recent advances in RAG have progressed from static retrieve-then-generate pipelines to more dynamic and adaptive frameworks (Wang et al., 2024b; Fan et al., 2024). For example, Self-RAG introduces ”reflection tokens” that enable the model to decide on-demand whether retrieval is necessary, avoiding the inefficiency of indiscriminate retrieval for every query (Asai et al., 2024). Similarly, DRAGIN (Su et al., 2024) proposes a dynamic retrieval mechanism that determines both when and what to retrieve during generation, further improving retrieval scheduling over static pipelines. In a similar vein, Active Retrieval Augmented Generation adopts an iterative retrieval process throughout generation, allowing the model to gather information as needed. Meanwhile, R²AG focuses on bridging the semantic gap between the retriever and the generator by leveraging more nuanced retrieval features (Ye et al., 2024). Despite these innovations, the practical deployment of RAG in commercial systems such as Perplexity AI and Google AI Overviews continues to reveal challenges in maintaining content quality and robust contextual grounding (Williams, 2024). Particularly, most existing RAG frameworks (static or dynamic) primarily assume that all required knowledge can be retrieved from pre-indexed textual corpora. Recent studies on inference-time scaling further show that allocating more test-time compute (e.g., iterative prompting or expanded retrieval) can improve answer quality in long-context settings. However, these approaches remain fundamentally limited to document retrieval over static indices and do not address queries whose answers must be generated through live interaction with external systems such as APIs or databases. In this work, TURA targets this complementary regime by enabling tool execution over dynamic, non-web, and transactional data sources while simultaneously optimizing latency through DAG-parallel planning and a distilled executor.
@@ 2.2. Tool-Augmented Agents
The foundational ReAct framework established a paradigm of interleaving reasoning and acting, where the model generates both thought processes and subsequent actions to interact with its environment. This synergy was shown to significantly improve task performance and interpretability. Building on this, Toolformer demonstrated how models could teach themselves to use tools in a self-supervised manner (Schick et al., 2023). The scale of tool integration was dramatically expanded by ToolLLM, which enabled models to leverage thousands of real-world APIs through the use of depth-first search-based decision trees (Qin et al., 2023). Recent work has also focused on optimizing the efficiency of tool use. LLMCompiler, for instance, introduced a Directed Acyclic Graph (DAG)-based approach for parallel tool calling, achieving a 3.6x speedup (Kim et al., 2024). Agent Q further improved success rates by combining Monte Carlo Tree Search with a self-critique mechanism (Putta et al., 2024). Despite these advances, current systems face critical limitations: (1) their workflows are often static and cannot adapt to the complexity of a given query; (2) they struggle with the semantic integration of heterogeneous tools and information sources; and (3) they suffer from coordination inefficiencies, as RAG and tool-augmented systems typically operate in isolation. In this work, TURA is designed to address these issues by proposing a unified architecture that dynamically coordinates retrieval decisions, tool selection, and response generation based on the specific characteristics and intent of the user’s query.
@@ 3. Problem Definition
Given a user’s natural language query $q$, we consider a collection of $N$ heterogeneous MCP servers, $M=\{M_{1},M_{2},\dots,M_{N}\}$, each encapsulating a specialized capability that is external to the language model. These capabilities span both static information access (e.g., indexed web search over documents) and dynamic information generation (e.g., transactional APIs or database queries).
The core problem is to answer the query $q$ by dynamically composing the functionalities of these MCP servers. Concretely, the system must autonomously decide whether to retrieve static web evidence or to execute dynamic tools, and coordinate their execution to produce a final synthesized answer $A$. We refer to any concrete composition and ordering of MCP server calls as an execution strategy, denoted by $\pi$. This formulation emphasizes the unified nature of our setting. Specifically, traditional web retrieval is itself treated as an MCP server (i.e., a search server), while external APIs and databases are also represented as MCP servers. As a result, both static retrieval and dynamic tool execution are orchestrated within a single planning-and-execution pipeline and operate over a unified interface, rather than being handled by separate subsystems.
The effectiveness of an execution strategy is governed by a fundamental trade-off between answer quality and system latency. We aim to maximize the quality of the final answer, $Q(A)$, which depends on the chosen strategy, while simultaneously satisfying a strict latency constraint, $L(\pi)$. Our objective is therefore to identify the optimal strategy $\pi^{*}$ that yields the best possible answer within a given time budget:
(1) $\pi^{*}=\operatornamewithlimits{arg\,max}_{\pi}Q(A(\pi))\quad\text{subject to}\quad L(\pi)\leq\tau_{\text{max}},$
where $\tau_{\text{max}}$ denotes the maximum permissible end-to-end latency. This formulation captures the central challenge of AI search in real-world production settings: effectively leveraging heterogeneous external capabilities under stringent performance constraints.
Figure 2. TURA Framework Overview. The framework consists of three stages: Intent-Aware MCP Server Retrieval, DAG-based Task Planner, and Distilled Agent Executor. The example shows the processing progress of TURA for the input query “Beijing trip June 10-15. Need hotel, 2-3 attractions and things to do”.
@@
4. Methodology
In this section, we first present the overall design of TURA, and then introduce the technical details of TURA.
4.1. Framework Overview of TURA
Fig. 2 illustrates the proposed framework TURA, which consists of three key modules: Intent-Aware MCP Server Retrieval Module, DAG-based Task Planner Module, and Distilled Agent Executor Module. Specifically, TURA (1) decomposes complex user queries into atomic sub-intents and performs dense retrieval over a semantically-indexed catalogue of available tools, encapsulating both static document collections and dynamic APIs, to identify relevant candidates for each sub-intent. Then, TURA constructs optimal and parallelizable execution plans by modeling sub-tasks and their data dependencies as a DAG, enabling the orchestration of complex, multi-hop reasoning chains across multiple tool calls. Eventually, to address critical inference latency barriers in production settings, TURA leverages a lightweight yet highly capable agent fine-tuned on curated expert trajectories using a novel mixed-rationale distillation technique to achieve comparable fidelity to proprietary LLMs at a fraction of the computational cost and latency.
@@ 4.2.1. LLM-based Multi-Intent Query Decomposition
User queries are often underspecified and multifaceted. A monolithic query may contain multiple, logically distinct information needs. To handle this, we employ a powerful LLM, $f_{\text{LLM-de}}$, configured with a specific prompt (shown in Appendix A.2) to function as a query decomposer. The LLM is instructed to parse the raw query $q$ and transform it into a structured set of atomic sub-queries $\mathcal{SQ}=\{sq_{1},sq_{2},\dots,sq_{k}\}$, which could be formulated as
(2) $\mathcal{SQ}=f_{\text{LLM-de}}(q).$
Each sub-query $sq_{j}\in\mathcal{SQ}$ is designed to be unambiguous and correspond to a single semantic intent (e.g., “find the city where the Forbidden City is located.”, “get weather forecast for a given city”). This decomposition transforms an ambiguous problem into a set of well-defined, tractable sub-problems. This strategy is inspired by recent advances in task decomposition for complex reasoning (Zhou et al., 2022).